OpenStudy (anonymous):

Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. y = x^2 + 3x + 2 y = x^2 + 2x + 1

OpenStudy (precal):

try factoring each one, the location where the function crosses the x axis is your answers

OpenStudy (precal):

I believe they are 2 solutions

OpenStudy (anonymous):

you can also solve this by using the quadratic formula \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (precal):

yes either method with work in this case, since these quadratics are factorable.

OpenStudy (anonymous):

y = y x^2+3x + 2 = x^2 + 2x + 1 3x + 2 = 2x + 1 x = -1 y = 0 So they intersect at the point(-1,0)

OpenStudy (anonymous):

is that right?

OpenStudy (precal):

is this a system of equations?

OpenStudy (anonymous):

on what they have in common

OpenStudy (anonymous):